OpenStudy (anonymous):

How many solutions are there to the following system of equations? 5x + 3y = 7 2x – y = 2 0 1 2 infinitely many

OpenStudy (anonymous):

1 use the coefficient property //

OpenStudy (anonymous):

How many solutions are there to the following system of equations? 6x + 12y = 11 4x + 8y = 3 (Points : 4) 0 1 2 infinitely many

OpenStudy (anonymous):

no real solution ..

OpenStudy (anonymous):

so it would be 0 right

OpenStudy (anonymous):

Solve the following system of equations by linear combination: 2d – e = 7 d + e = 5 (Points : 4) There is no solution. There are an infinite number of solutions. The solution is (3, –1). The solution is (4, 1).

OpenStudy (anonymous):

add both the eqns.. and get d=4 put in anyone obtain e